OpenStudy (anonymous):

Choose the correct simplification of x^2/x^5. A. 1/x^3 B. x^3 C. x^7 D. 1/x^7 is it B?

Directrix (directrix):

A. 1/x^3

Directrix (directrix):

Not B because x^2 / x^5 = x * x / x * x* x * x * x = x*x ---> site is about to shut down -- will finish this later today

OpenStudy (anonymous):

Thanks and okay, I be here! :)

Directrix (directrix):

You're here

Directrix (directrix):

Maybe I can finish and if this site restarts or whatever, then let it be . It crashed big time last night with no notice.

OpenStudy (anonymous):

Idk.. But if your willing to, I'm going to post 13 questions on yahoo answers. I just need them to be checked, I've already done them all I just want to make sure there correct?

OpenStudy (anonymous):

whoa the website is bugging out, I didn't send that last reply. hahaha

Directrix (directrix):

is it B? ====== Not B because x^2 / x^5 = x * x / x * x* x * x * x = [x/x * x/x] / x * x * x = 1 * 1 / x * x * x = 1/x^3. The extra x terms are in the denominator. Your answer goes with the following problem: x^5/ x^2

Directrix (directrix):

I have worked on Yahoo! Answers. It's an interesting place.

Directrix (directrix):

Finding your questions on Yahoo! Answers might be dicey.

OpenStudy (anonymous):

This was posted before I got the right answer from you on question 10 but this is the link to it. http://answers.yahoo.com/question/index?qid=20120224121349AAHV1UZ

Directrix (directrix):

Look what popped up over there in the land of the Yahoos. This question has been deleted Questions on Yahoo! Answers are sometimes deleted according to our Community Guidelines.

OpenStudy (anonymous):

ohhh no wait one sec

OpenStudy (anonymous):

Try this one: http://answers.yahoo.com/question/index?qid=20120224121349AAHV1UZ

OpenStudy (anonymous):

They delete people questions sometimes because they cheat, but I'm just checking my answers so idk why mine was deleted??

OpenStudy (anonymous):

Did you find it?

OpenStudy (anonymous):

I rechecked them and I did get number 2 wrong, its A instead of B. But other than that, I think there all correct.

OpenStudy (anonymous):

Thanks for all of your help!! :)

Directrix (directrix):

No, the link went to a page that said the posting had been deleted.

OpenStudy (anonymous):

hmmm, idk why. It's working for me. You can go to a cached version?

Directrix (directrix):

I found it. I have questions about problem #2, #8, #10, and #11. Perhaps you left them incorrect for the Yahoo checker.

OpenStudy (anonymous):

Yeah, 2 is A, 8 is D, 10 is a and 11 is A. I found that out after I had already posted it :)